OpenStudy (anonymous):

The range of F(x) = 5cosx is _____. -5 ≤ F(x) ≤ 5 -5 ≤ F(x) ≤ 0 0 ≤ F(x) ≤ 5

OpenStudy (solomonzelman):

the range of cos(x) is \(\large\color{black}{ [-1,1] }\), the 5 stretches it by 5 both ways, so it will be, \(\large\color{black}{ [-5,5] }\) .

OpenStudy (solomonzelman):

I'll get a graph to chec, in a sec.

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/wwiklpzygo

OpenStudy (michele_laino):

I think the first option is your answer, because the natural limits of the function cos(x), are: \[-1\le \cos x \le +1\] then we conclude that: \[-5\le 5*\cos (x)\le +5\]

OpenStudy (solomonzelman):

Yes;)

OpenStudy (anonymous):

thank you both

OpenStudy (michele_laino):

@mudwagaman9 thank you!